McCallister's just purchased $16,600 worth of inventory. The terms of the sale were 2/10, net 40. What is the implicit interest?
Answer #1
Discount to be withdrawn= 2 (i.e. interest )
On amount to be paid = 100-2 = 98
period of additional credit allowed = 40-10 = 30
Implicit rate of interest = 2/ 98 * 365/30 =24.83%
Answer is 24.83%